An economy produces three goods: orange, apple and orange juice. It is assumed that half of the orange are bought and consumed as final good; the other half is used to produce orange juice. In 2019, 200 KGs of oranges are produced at Rs.40/kg, 500 KGs of apples are produced at 120/kg, and 200 Glasses of orange juice are produced at Rs.150/glass. In 2020, 220 KGs of oranges are produced at Rs.50/kg, 450 KGs of apples are produced at 140/kg, and 230 glasses of orange juice are produced at Rs.135/glass. Assuming 2019 is the base year; calculate the nominal GDP in 2020


1
Expert's answer
2021-05-04T11:18:26-0400

Nominal GDP = sum of the values of all the final goods

For year 2020;

Out of 220 kgs of oranges, 110 kgs are consumed as oranges and 110kgs are consumed for making orange juice

Hence, final goods = 110kgs of oranges valued at Rs 50/kg

Final amount of apples = 450kgs at Rs 140/kg

Final amount of orange juice = 230 glasses at Rs 135/glass

 

Nominal GDP"= (110\times 50) + (450\times 140) + (230\times135) = Rs 99,550"